# SPADE: Structured Pruning and Adaptive Distillation for Efficient LLM-TTS

### v2

@@ 1 Introduction
Fig. 2: WLI and cosine-based layer importance of (a) CosyVoice 2 and (b) LLaSA. High WLI indicates that WER increases significantly when the layer is removed, and high cosine-based importance indicates the input and output latents of the layer are dissimilar. We found that, based on WLI, the layers in the beginning, middle, and the end contributes critically to the performance. Our method prunes the model by removing layers with least contribution to the performance.
@@ 2 Proposed Method
Our framework first explores the importance of each layer in LLM backbone and prune non-essential layers to compress the model. Subsequently, an efficient knowledge distillation is applied to effectively restore the performance of the pruned model.
@@ 2.1 Model compression through pruning
where ${\theta}_{\setminus i}$ denotes the model parameters without the $i$-th layer, $\mathcal{D}$ is a subset of the evaluation set, and $({\bm{x}}_{1},{\bm{y}}_{1})$ and $({\bm{x}}_{2},{\bm{y}}_{2})$ denote reference and query text-audio pairs, respectively. Specifically, a layer is considered important only if its absence causes significant degradation in WER. Unlike cosine-based layer importance, which only estimates the difference between inputs and outputs of within a layer, WLI directly measures the contribution of each layer to the final performance. As shown in Fig 2 evaluated WLI with Whisper [25], and found that many of the layers have negligible impact on performance, indicating a significant redundancy and aligning with the findings in text-domain LLMs [13]. Moreover, we find that the earliest, central, and final layers consistently emerge as important across different LLM-TTS models. Based on the analysis, we prune transformer layers with low WLI values from the LLM backbones.
@@ 2.2 Recovering original performance
Here, the Cross-Entropy loss $\mathcal{L}_{CE}$ is responsible for the supervised component that directly guides the output distribution. The distillation component comprises 4 elements: Embedding reconstruction loss $\mathcal{L}_{e}$, alignment losses on logit $\mathcal{L}_{logit}$, latent $\mathcal{L}_{l}$, and attention $\mathcal{L}_{a}$, following [24]. To provide more stability, we implement the $\mathcal{L}_{logit}$ with mixed distribution by leveraging Skew KL Divergence [16]. For $\mathcal{L}_{e},\mathcal{L}_{l}$, and $\mathcal{L}_{a}$, we calculate Mean Squared Error (MSE) of embedding outputs, intermediate latents, and attention matrices between the teacher and student, respectively. To maximize the distillation of teacher’s knowledge, we propose to dynamically select layers to apply loss for $\mathcal{L}_{l}$ and $\mathcal{L}_{a}$. Specifically, as shown in Fig 1, the target values (latent, attention map) for the student layer ($l_{n}$) are derived from the last layer before the next retained layer ($l_{m+2}$) from the teacher model. While simple, this approach allows the pruned student model to retain not only its original capability but also those of the removed layers, culminating in a smaller yet more compact model. Finally, we combine the supervised and distillation components by adjustable weight alphaalpha to balance the influence of each, where the value is empirically set to $0.25$ to deliberately provide stronger supervised guidance.
@@ 4.1 Effect on performance after pruning
Table 1 presents the main results of applying SPADE to CosyVoice 2 and LLaSA with different configurations. For CosyVoice, pruning to 12 layers halves the depth, reduces parameters by $39.7\%$ and accelerates inference by $42.6\%$. Moreover, effective VRAM usage is reduced by $14\%$, as depicted in Fig. 3. Notably, these gains come with no significant degradation in both quantitative and qualitative metrics, showing slight increase of 0.68 in WER for challenging Seed-TTS dataset and 0.11 reduction in NMOS. A more aggressive variant with only 9 layers further reduces parameters by $49.2\%$, improves RTF by $45.9\%$, and lowers VRAM usage by $17\%$. While this extreme setting renders additional increase in WER, perceptual metrics such as NMOS, SS, and UTMOS remain stable, suggesting that SPADE enables flexible trade-offs between efficiency and intelligibility. To confirm that our framework generalizes beyond a single backbone, we conduct experiment on LLaSA, a larger variant based on speech codec. Here, pruning removes half the layers, decreases parameters by $23.5\%$, improves RTF by $29.3\%$ ($1.41\times$ speed-up), and reduces VRAM by $20\%$. The result suggests that, while most of the metrics lie in acceptable range, the performance degradation is relatively larger compared to CosyVoice 2. Based on the analysis in Fig 2444Metrics are reported based on official checkpoints, we expect it attributes to overall high WLI values across all layers, indicating each layer contributes similarly to the performance.
@@ 4.2 Ablation study
We evaluate the effectiveness of the proposed methodology with systematic ablation study. Table 2 presents experiments on CosyVoice 2 with LibriTTS test-clean. First, the proposed WLI-based pruning is replaced with existing cosine-based pruning. The result shows a notable increase in WER and CER, demonstrating the pruning based on WLI, a metric directly related to intelligibility, successfully prevents performance degradation. Moreover, dynamic distillation loss is removed and student layer is distilled only with the information from the corresponding teacher layer. The overall decrease in performance suggests that adaptively choosing the target in distillation training is simple yet plays a significant role.

### v3

@@ 1 Introduction
Fig. 2: WLI and cosine-based layer importance of (a) CosyVoice 2 and (b) LLaSA. High WLI indicates that WER increases significantly when the layer is removed, and high cosine-based importance indicates the input and output latents of the layer are dissimilar. We found that, based on WLI, the layers in the beginning, middle, and the end contribute critically to performance. Our method prunes the model by removing layers with least contribution to the performance.
@@ 2 Proposed Method
Our framework first explores the importance of each layer in LLM backbone and prunes non-essential layers to compress the model. Subsequently, an efficient knowledge distillation is applied to effectively restore the performance of the pruned model.
@@ 2.1 Model compression through pruning
where ${{\bm{\theta}}}_{\setminus i}$ denotes the model parameters without the $i$-th layer, $\mathcal{D}$ is a subset of the evaluation set, and $({\bm{x}}_{1},{\bm{y}}_{1})$ and $({\bm{x}}_{2},{\bm{y}}_{2})$ denote reference and query text-audio pairs, respectively. Specifically, a layer is considered important only if its absence causes significant degradation in WER. Unlike cosine-based layer importance, which only estimates the difference between inputs and outputs within a layer, WLI directly measures the contribution of each layer to the final performance. As shown in Fig. 2 evaluated WLI with Whisper [19], and found that many of the layers have negligible impact on performance, indicating a significant redundancy and aligning with the findings in text-domain LLMs [8]. Moreover, we find that the earliest, central, and final layers consistently emerge as important across different LLM-TTS models. Based on the analysis, we prune transformer layers with low WLI values from the LLM backbones.
@@ 2.2 Recovering original performance
Here, the Cross-Entropy loss $\mathcal{L}_{CE}$ is responsible for the supervised component that directly guides the output distribution. The distillation component comprises 4 elements: Embedding reconstruction loss $\mathcal{L}_{e}$, alignment losses on logit $\mathcal{L}_{logit}$, latent $\mathcal{L}_{l}$, and attention $\mathcal{L}_{a}$, following [17]. To provide more stability, we implement the $\mathcal{L}_{logit}$ with mixed distribution by leveraging Skew KL Divergence [14]. For $\mathcal{L}_{e},\mathcal{L}_{l}$, and $\mathcal{L}_{a}$, we calculate Mean Squared Error (MSE) of embedding outputs, intermediate latents, and attention matrices between the teacher and student, respectively. To maximize the distillation of teacher’s knowledge, we propose to dynamically select layers to apply loss for $\mathcal{L}_{l}$ and $\mathcal{L}_{a}$. Specifically, as shown in Fig. 1, the target values (latent, attention map) for the student layer ($l_{n}$) are derived from the last layer before the next retained layer ($l_{m+2}$) from the teacher model. While simple, this approach allows the pruned student model to retain not only its original capability but also those of the removed layers, culminating in a smaller yet more compact model. Finally, we combine the supervised and distillation components by adjustable weight alphaalpha to balance the influence of each, where the value is empirically set to $0.25$ to deliberately provide stronger supervised guidance.
@@ 4.1 Effect on performance after pruning
Table 1 presents the main results of applying SPADE to CosyVoice 2 and LLaSA with different configurations. For CosyVoice, pruning to 12 layers halves the depth, reduces parameters by $39.7\%$ and accelerates inference by $42.6\%$. Moreover, effective VRAM usage is reduced by $14\%$, as depicted in Fig. 3. In particular, these gains come with no significant degradation in both quantitative and qualitative metrics, showing a slight increase of 0.68 in WER for the challenging Seed-TTS dataset and a 0.11 reduction in NMOS. A more aggressive variant with only 9 layers further reduces parameters by $49.2\%$, improves RTF by $45.9\%$, and lowers VRAM usage by $17\%$. While this extreme setting renders additional increase in WER, perceptual metrics such as NMOS, SS, and UTMOS remain stable, suggesting that SPADE enables flexible trade-offs between efficiency and intelligibility. To confirm that our framework generalizes beyond a single backbone, we conduct experiment on LLaSA, a larger variant based on speech codec. Here, pruning removes half the layers, decreases parameters by $23.5\%$, improves RTF by $29.3\%$ ($1.41\times$ speed-up), and reduces VRAM by $20\%$. The result suggests that, while most of the metrics lie in acceptable range, the performance degradation is relatively larger compared to CosyVoice 2. Based on the analysis in Fig. 2444Metrics are reported based on official checkpoints, we expect it attributes to overall high WLI values across all layers, indicating each layer contributes similarly to the performance.
@@ 4.2 Ablation study
We evaluate the effectiveness of the proposed methodology with a systematic ablation study. Table 2 presents experiments on CosyVoice 2 with LibriTTS test-clean. First, the proposed WLI-based pruning is replaced with existing cosine-based pruning. The result shows a notable increase in WER and CER, demonstrating the pruning based on WLI, a metric directly related to intelligibility, successfully prevents performance degradation. Moreover, dynamic distillation loss is removed and student layer is distilled only with the information from the corresponding teacher layer. The overall decrease in performance suggests that adaptively choosing the target in distillation training is simple yet plays a significant role.
@@ 5 Conclusion
6 acknowledgement
This work was partially supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT, RS-2025-02263977, Development of Communication Platform supporting User Anonymization and Finger Spelling-Based Input Interface for Protecting the Privacy of Deaf Individuals).